# DMODE: Differential Monocular Object Distance Estimation Module without Class Specific Information

### v2

@@ I INTRODUCTION
The accuracy of class-specific object detection relies on matching the training environment. For example, in a test scenario involving toy objects, a toy car at the same distance as a real car will appear much smaller but may still be detected as a “Car” by object classification networks. Similarly, when an object is presented in the camera field of view while tilted, a class-specific approach can only detect the distance correctly if the object is in the dataset with the exact pose, which is unlikely or requires an enormous dataset. Finally, the precision in class-specific approaches with multiple classes is limited to the accuracy of the classification technique. These limitations don’t affect class-agnostic approaches [13], which don’t require knowledge of expected object sizes at varying distances.
@@ II-A Monocular depth estimation
Depth estimation has been approached using DNNs, such as continuous condition random fields for image patches proposed by [14]. The accuracy of depth estimation was improved by [15], who incorporated ordinal regression into the depth estimation network and used scale-increasing discretization to convert continuous depth data into an ordinal class vector. These techniques require significant manpower and computing resources, as well as specific training images and corresponding depth maps for each pixel [16]. In the absence of a depth image serving as ground truth (GT), unsupervised training might utilize additional depth cues from stereo images [17] or optical flows [18]. [19] introduced a novel deep visual-inertial odometry and depth estimation framework to enhance the precision of depth estimation and ego-motion using image sequences and inertial measurement unit (IMU) raw data. However, unsupervised depth estimation methods have inherent scale ambiguity and poor performance due to the lack of perfect GT and geometric constraints [19].
@@ IV-A Theoretical analysis
[…]
We assume objects moving with constant velocity which requires $q=2$. Hence, $\Delta D_{1}=\Delta D_{2}=\Delta D$ and Eq. (9) will be:
@@ IV-B Network architecture
As shown in Fig. 2, our network is a sequential architecture consisting of a 2D convolutional neural network (CNN) and three multi-layer perceptrons (MLPs) with batch normalization implemented between each layer. After processing the detected object from OTN ($\mathcal{F}_{OTN}$), three consecutive frames are stacked and passed through a mapping function ($\mathcal{F}_{MAP}$) comprised of an MLP with a soft-plus activation function. As shown in the upper right corner of Fig. 2, upon the availability of 3 frames of an object, we resize and crop them to the size of $224\times 224$. We employ ResNet-18 ($\mathcal{F}_{ResNet18}$) [29] by inputting selected bounding box (BBox) information to yield $1000$ parameters as latent variables ($\mathcal{Z}$).
@@ IV-C Learning rules
where, $\phi^{*}$ represents the corresponding GT of Cartesian space, and $c=0.2.\max{(|\phi-\phi^{*}|)}$. The BerHu loss function integrates $L_{1}$ and $L_{2}$ features and adapts according to the $c$ value. Contrary to the $L_{2}$ loss function, the BerHu loss function converges more rapidly while effectively preserving small residuals, as it is differentiable at the point $c$ where the transition between $L_{1}$ and $L_{2}$ losses takes place. This enables the network to leverage the shrinkage power of $L_{1}$ norm, thereby avoiding low gradients for relatively small residuals. The final loss is defined as follows:
@@ VII Conclusion
We introduced a class-agnostic method for object distance estimation, leveraging the object’s evolving appearance and camera motion. This property enables TL with target class samples, yielding satisfactory multi-class distance estimation results. However, our approach has two main limitation: a time delay is required before distance estimation, and accurate object tracking over three key frames, matching the training intervals, is essential. Moreover, The network architecture currently employs three input channels, suitable for the basic version assuming constant object velocity. While results are generally good, outliers may occur for rapidly changing objects. Future work may involve increasing the channel capacity of our 2D network to improve performance in such cases.

### v3

@@ I INTRODUCTION
The accuracy of class-specific object detection relies on matching the training environment [13]. For example, in a test scenario involving toy objects, a toy car at the same distance as a real car will appear much smaller but may still be detected as a “Car” by object classification networks. Similarly, when an object is presented in the camera field of view while tilted, a class-specific approach can only detect the distance correctly if the object is in the dataset with the exact pose, which is unlikely or requires an enormous dataset. Finally, the precision in class-specific approaches with multiple classes is limited to the accuracy of the classification technique. These limitations don’t affect class-agnostic approaches [14], which don’t require knowledge of expected object sizes at varying distances.
@@ II-A Monocular depth estimation
Depth estimation has been approached using DNNs, such as continuous condition random fields for image patches proposed by [15]. The accuracy of depth estimation was improved by [16], who incorporated ordinal regression into the depth estimation network and used scale-increasing discretization to convert continuous depth data into an ordinal class vector. These techniques require significant manpower and computing resources, as well as specific training images and corresponding depth maps for each pixel [17]. In the absence of a depth image serving as ground truth (GT), unsupervised training might utilize additional depth cues from stereo images [18] or optical flows [19]. Authors in [20] introduced a novel deep visual-inertial odometry and depth estimation framework to enhance the precision of depth estimation and ego-motion using image sequences and inertial measurement unit (IMU) raw data. However, unsupervised depth estimation methods have inherent scale ambiguity and poor performance due to the lack of perfect GT and geometric constraints [20].
@@ IV-A Theoretical analysis
$(1-p_{1})d_{1}=\Delta C_{1},~{}~{}\textit{if}~{}q=1$ (11)
By assuming objects moving with constant velocity ($q=2$), we have $\Delta D_{1}=\Delta D_{2}=\Delta D$ and Eq. (9) is updated as:
@@ IV-B Network architecture
As shown in Fig. 2, the sequential network, inspired from [30], consist of a 2D convolutional neural network (CNN) and three multi-layer perceptrons (MLPs) with batch normalization implemented between each layer. After processing the detected object from OTN ($\mathcal{F}_{OTN}$), three consecutive frames are stacked and passed through a mapping function ($\mathcal{F}_{MAP}$) comprised of an MLP with a soft-plus activation function. As shown in the upper right corner of Fig. 2, upon the availability of 3 frames of an object, we resize and crop them to the size of $224\times 224$. We employ ResNet-18 ($\mathcal{F}_{ResNet18}$) [31] by inputting selected bounding box (BBox) information to yield $1000$ parameters as latent variables ($\mathcal{Z}$).
@@ IV-C Learning rules
where, $\phi^{*}$ represents the corresponding GT of Cartesian space, and $c=0.2.\max{(|\phi-\phi^{*}|)}$. The BerHu loss function integrates $L_{1}$ and $L_{2}$ features and adapts according to the $c$ value. Contrary to the $L_{2}$ loss function, the BerHu loss function converges more rapidly while effectively preserving small residuals, as it is differentiable at the point $c$ where the transition between $L_{1}$ and $L_{2}$ losses takes place. This enables the network to leverage the shrinkage power of $L_{1}$ norm, thereby avoiding low gradients for relatively small residuals. The final loss inspired from multi-task training in [33] is defined as follows:
@@ VII Conclusion
We introduced a class-agnostic method for object distance estimation, leveraging the object’s evolving appearance and camera motion. This property enables TL with target class samples, yielding satisfactory multi-class distance estimation. The limitation of our approach includes time delay and accurate object tracking over three key frames, matching the training intervals, is essential. Moreover, The network architecture currently employs three input channels, suitable for constant object velocity. While results are generally good, outliers may occur for rapidly changing objects. Future work may involve increasing the channel capacity of our 2D network to improve performance in such cases.